OpenStudy (anonymous):

The perimeter of an equilateral triangle is 12 feet. Find the area of the triangle. 8√3 8 4√3

OpenStudy (anonymous):

a=1/2bh or bh/2 the base and the height are equal to 12/3 =4 4*4=16 16/2=8 your answer is 8